OpenStudy (anonymous):

Rewrite without parenthesis and simplify . (y+6)^2

OpenStudy (mathmale):

y+6 is a binomial. So, your task is to square this binomial. Note that squaring y+6 produces the same result as multiplying (y+6)(y+6). Try using the FOIL method on the latter.

OpenStudy (anonymous):

There are these 3 identities that you should remember: (very handy if you do any algebra) (a+b)^2=a^2+2ab+b^2 (a-b)^2-a^2-2ab+b^2 (a+b)(a-b)=a^2-b^2

OpenStudy (anonymous):

\[y^2+12y+36\]

OpenStudy (anonymous):

Thank you guys . (: